HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

(SUB-) MILLIMETER WAVE SPECTRUM OF METHYLENE CHLORIDE, CH_{2}^{35}Cl_{2}, UP TO 1.1 THz

\underline{O. N. ULENIKOV} , O. V. GROMOVA, E. S. BEKHTEREVA, Research School of

High–Energy Physics, National Research Tomsk Polytechnic University, Tomsk 634050, Russia; H. S. P. MÜLLER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; L. MARGULÈS, UMR8523–PhLAM, Université de Lille,

59000 Lille, France

Methylene chloride is of fundamental interest as a small and fairly light molecule. In addition, it is an important industrial solvent, and its identification by spectroscopic means in industrial transportation and in the processing of petroleum products. We have recorded its high resolution rotational spectrum in two laboratories throughout most of the frequencies from 75 GHz to 1.09 THz. More than 4000 rotational transitions have been identified with J ≤ 108 and K_{a} ≤ 20. These data have been fit to within 20 kHz employing spectroscopic parameters up to 10th order.

The study was supported in part by The Ministry of Science and High Education of the Russian Federation in part of the Science program (Project FSWW–2023– 0003).

p-number: p074  Submitted on Mon, 12 Jun 2023 08:46:22 +0200